You have $65,000.

You put 21​% of your money in a stock with an expected return of 15​%, $31,000 in a stock with an expected return of 13​%, and the rest in a stock with an expected return of 19​%.

What is the expected return of your​ portfolio?

The expected return of your portfolio is? (Round to two decimal​ places.)

Please show steps taken for all questions.
Answer #1

w2 = $31,000/$65,000 = 0.4769

w3 = 1 - w1 - w2 = 1 - 21% - 47.69% = 31.31%

Expected return for portfolio = 21% * 15% + 47.69% * 13% + 31.31% * 19%

Expected return of portfolio = 3.15% + 6.20% + 5.95%

Expected return of portfolio = 15.30%